OpenStudy (frostbite):

A function is given by: Y=(x+3x^2+3x^3+10x^4)/(1+4x+x^2+4x^3+10x^4 ),x≥0 Explain and argue for that Y/x assume a global maximum for x>0 (have done that). And determine the maximum with 3 decimals (just setting a calculator to get a lot of digits or zoom very close to the graph is not acceptable).

OpenStudy (frostbite):

I was thinking about determine the decimals by using the Newton/Raphson-method

OpenStudy (anonymous):

did you try to graph this one? so that you can see the MAX so you said the Max is Y/X ?

OpenStudy (anonymous):

Y/x=(1+3x+3x^2+10x^3)/(1+4x+x^2+4x^3+10x^4 )

OpenStudy (anonymous):

there are local min and local max

OpenStudy (frostbite):

Yea beside we may not plot the function

OpenStudy (anonymous):

to find local max n min find the derivative and equate it to zero to find the x

OpenStudy (frostbite):

I add another comment that also is stated in the problem: "You must explain your decimals reliability"

OpenStudy (anonymous):

set dy/dx= (-100x^6-60x^5-92x^4+16x^3+27x^2+4x-1)/(10x^4+4x^3+x2+4x+1)^2=0 x=0.131 x=0.512

OpenStudy (anonymous):

try also the Newton/Raphson-method and compare

OpenStudy (frostbite):

I know that you very esay can find the values but just useing a caculator which i have done, but the whole idea is that we are going to estamate a aproximated number without useing electronic help

OpenStudy (frostbite):

by using*

OpenStudy (anonymous):

do ypu have a graphing calculator? can you use that?

OpenStudy (frostbite):

No, we not use a graph either.

OpenStudy (anonymous):

(just setting a calculator to get a lot of digits or zoom very close to the graph is not acceptable).--->>> isnt this means that graphic calcu is ok that you cant zoom it only....

OpenStudy (anonymous):

well if you cant use a grapher try the Newton/Raphson-method

OpenStudy (frostbite):

And that is what i can't not very good useing it.

OpenStudy (anonymous):

what the grapher can do for us is to help us find the answer and compare what you have done on your solutions..lol,,....well good luck now